OpenStudy (anonymous):

If the frequency of a particle in SHM is f ,then how frequency of variation of K.E is 2f???

OpenStudy (anonymous):

let for a partical to complete it's 1 cycle it takes time T but as we know that KE has no direction it's value will be changed in half time so frequency will be doubled for it it increases form 0 to max then max to 0 in T/2 times so f'=(1/(T/2))=2/T=2f